If the side of a triangle are 7,6,and 9 the triangle is an
Found 2 solutions by unlockmath, kkasko:
Answer by unlockmath(1688)

Hello,
It's an isosceles triangle.
Use this formula:
A^2+B^2=C^2
7^2+6^2=9^2
49+36=81
85 doesn't equal 81 so we know it's not a right triangle and two sides are not equal so it only leaves us with an isosceles triangle.
Make sense?
RJ
www.math-unlock.com

Answer by kkasko(46)

If the side of a triangle are 7,6,and 9 the triangle is an scalene triangle since no sides are equal.